### v1

# Conditioning Diffusions Using Malliavin Calculus

Abstract
In stochastic optimal control and conditional generative modelling, a central computational task is to modify a reference diffusion process to maximise a given terminal-time reward. Most existing methods require this reward to be differentiable, using gradients to steer the diffusion towards favourable outcomes. However, in many practical settings, like diffusion bridges, the reward is singular, taking an infinite value if the target is hit and zero otherwise. We introduce a novel framework, based on Malliavin calculus and path-space integration by parts, that enables the development of methods robust to such singular rewards. This allows our approach to handle a broad range of applications, including classification, diffusion bridges, and conditioning without the need for artificial observational noise. We demonstrate that our approach offers stable and reliable training, outperforming existing techniques.
Bridges, h-transform, conditioning, diffusions, SDEs
1 Introduction
Simulating conditioned diffusions is a central computational task in many applications, ranging from molecular dynamics and physical chemistry (Dellago et al., 2002; Bolhuis et al., 2002; Vanden-Eijnden et al., 2010) and genetics (Wang et al., 2011) to finance, econometrics (Bladt & Sørensen, 2014; Elerian et al., 2001; Durham & Gallant, 2002), generative modelling (Dhariwal & Nichol, 2021; Ho & Salimans, 2022) and evolutionary biology (Arnaudon et al., 2023, 2017; Baker et al., 2024).
Reference System. To demonstrate some of the core ideas, let us consider a diffusion process of the form
$\mathrm{d}X_{t}=b(X_{t})\,\mathrm{d}t+\mathrm{d}B_{t},$ (1)
where the drift vector field $b$ is given–either from a learned generative model or from a physically or financially motivated system. Samples from (1) can be obtained by straightforward numerical simulation.
Conditioning on Observations. However, in many settings, we also have an observation
$Y=G(X_{T})$ (2)
and would like to condition (1) on the event $Y=y$ (which could fix a property of the sample, or an area of the state space). If $Y=y$ is a rare event, then simulating many paths of (1) until $y$ occurs can be very wasteful.
Instead, we augment this setup by introducing
$\mathrm{d}X_{t}=b(X_{t})\,\mathrm{d}t+u_{t}(X_{t})\,\mathrm{d}t+\mathrm{d}B_{t},$ (3)
so that a suitable control $u_{t}$ can redirect the trajectory in light of new data or particular constraints (e.g. conditioning on rare events or outputting a sample with a specific property).
Figure 1: The figure depicts particle trajectories in a double well potential, with the background color indicating potential intensity. The potential has two metastable states at $x=1$ and $x=-1$. In (a) we observe that under the diffusion dynamics, particles initialised at $x=1$ typically remain confined to their well, rarely crossing the barrier to $x=-1$. On the right, in (b), the diffusion is conditioned on the rare event of transitioning between the two metastable states.
Likelihoods and Rewards with Gradients. The function $G$ induces a likelihood or reward
$g(x;y):=p(Y=y|X_{T}=x).$ (4)
Current methods, framed in terms of stochastic optimal control (Domingo-Enrich, 2024; Zhang & Chen, 2022; Berner et al., 2024), critically rely on gradient information $\nabla g$ or $\nabla\log g$ to guide the process and learn $u_{t}$.
Singular Rewards. In many settings there is no gradient information available for $g$. One case where this happens is if $G$ is an indicator function or discontinuous as it would be the case for classification. But even if $G$ is smooth but deterministic, the induced likelihood $g$ is often singular if one does not add artifical noise to the observations. Even in the seemingly straightforward case of $G=\text{Id}$, the reward would be given by a Dirac delta distribution $g(x_{T};y)=\delta_{y}(x_{T})$, which is not even continuous. The case of $G=\text{Id}$ conditions diffusions to end at a specific state and is known under the term diffusion bridges. Due to $g$ being a Dirac it is in some sense the most challenging setting and will be a guiding problem for us in the development of the methodology.
Integration by Parts on Path Space. In this paper, we circumvent these issues altogether by constructing numerical methods that remain unaffected by singularities in the likelihood $g$. The key idea is best introduced via an analogy: replace, for the moment, the trajectory space associated with 1 by the real line. The classical integration-by-parts identity
$\int_{-\infty}^{\infty}\partial_{x}g(x)f(x)\,\mathrm{d}x=-\int_{-\infty}^{\infty}g(x)\partial_{x}f(x)\,\mathrm{d}x,$ (5)
valid provided $f$ and $g$ vanish sufficiently fast at infinity, offers a blueprint for handling singularities: the left-hand side can be given a rigorous meaning even if $g$ lacks differentiability, as long as $f$ is smooth. Numerically, large (exploding) gradients can be avoided by shifting differentiation onto the factor with better properties.
Outline and Contributions. Guided by 5, we develop loss functions $\mathcal{L}(u)$ whose unique minimisers $u_{t}(x;y)$ drive the controlled diffusion in 3 to form bridges of 1 for any condition $Y=y$. Towards this goal, we
  • recall the well-known (Eberle, 2015; Denker et al., 2024; Shi et al., 2024; Du et al., 2024) connection of diffusion bridges to Doob’s $h$-transform in Section 2.1, in particular the relevance of conditional score functions,
  • lift 5 to integration by parts on the space of trajectories; the role of Lebesgue measure $\mathrm{d}x$ is replaced by the law of $(X_{t})_{0\leq t\leq T}$, and the ordinary derivative $\partial_{x}$ is replaced by the Malliavin derivative (Nualart, 2006). Based on this, we derive a novel formula for conditional scores that generalises Tweedie’s formula (Efron, 2011) for denoising score matching (Vincent, 2011),
  • discuss implementation details and showcase numerical performance in Section 4.
Furthermore, we recover existing methods for stochastic optimal control from our framework, see Section 3.2.
2 Theoretical Background and Main Result
Theorem 2.1 below serves as the mathematical foundation of our methodology. Before presenting it, we introduce the necessary notation and assumptions.
Throughout, we consider diffusion processes of the form
$\mathrm{d}X_{t}=b_{t}(X_{t})\,\mathrm{d}t+\sigma_{t}(X_{t})\,\mathrm{d}B_{t},\qquad X_{0}=x_{0},$ (6)
where $b:[0,T]\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a smooth drift of at most linear growth, $x_{0}\in\mathbb{R}^{n}$ is a fixed initial state, $B_{t}$ is a standard Brownian motion, and $\sigma:[0,T]\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$ specifies the volatility, assumed to be symmetric, strictly positive definite and bounded, with bounded inverse. A key component is the Jacobian $J_{t\mid s}$ associated to 6, which is a matrix-valued stochastic process satisfying
$\mathrm{d}J_{t\mid s}=\nabla b_{t}(X_{t})J_{t\mid s}\,\mathrm{d}t+\nabla\sigma_{t}(X_{t})J_{t\mid s}\,\mathrm{d}B_{t},$ (7)
with initial condition $J_{s\mid s}=\operatorname{Id}$, for fixed $s\in[0,T]$. Intuitively, $J_{t\mid s}$ measures the sensitivity of 6 at time $t$ with respect to a small perturbation at an earlier time $s<t$. More precisely, it represents the derivative process, i.e. $J_{t\mid s}=\nabla_{X_{s}}X_{t}$ (Williams & Rogers, 1979, Chapter V.13). The process $J_{t\mid s}$ plays a central role in adjoint methods for gradient computation (Li et al., 2020) and stochastic optimal control (Domingo-Enrich et al., 2024b), and we highlight the fact that simulating the full matrix-valued evolution is typically unnecessary as only matrix-vector products are required (see Appendix A). Our main theoretical contribution can now be stated as follows.
Theorem 2.1.
Let $\alpha:[0,T]\rightarrow\mathbb{R}^{n\times n}$ be a matrix-valued differentiable function such that $A_{T\mid s}\coloneqq\alpha_{T}-\alpha_{s}$ is invertible for all $s\in[0,T)$. Define the score process
$\mathcal{S}_{s}\coloneqq A^{-1}_{T\mid s}\int_{s}^{T}\alpha_{t}^{\prime}J_{t\mid s}^{\top}(\sigma_{t}(X_{t})^{\top})^{-1}\,\mathrm{d}B_{t},$ (8)
as well as the loss functional
$\mathcal{L}(u)=\mathbb{E}\left[\int_{0}^{T}\left\|u_{s}(X_{s};Y)-\mathcal{S}_{s}\right\|^{2}\,\mathrm{d}s\right],$ (9)
where the expectation is taken with respect to the injected noise $(B_{t})_{0\leq t\leq T}$ driving 6, 7 and 8, as well as potential noise in $G$ in 2. Then $\mathcal{L}$ admits a unique minimiser $u^{*}$, and for any $y\in\mathbb{R}^{d}$, the law of
$\mathrm{d}X_{t}=$ $~{}b_{t}(X_{t})\,\mathrm{d}t+\sigma_{t}(X_{t})\sigma_{t}(X_{t})^{\top}~{}u^{*}_{t}(X_{t};y)\,\mathrm{d}t$ (10)
$+\sigma_{t}(X_{t})\,\mathrm{d}B_{t}$
coincides with the conditional law of 6, given $Y=y$.
Based on Theorem 2.1, we propose to (i) estimate the expectation in 9 using Monte Carlo, (ii) parameterise the drift $u_{t}$ (including the conditioning) by a neural network, and (iii) learn the parameters of $u_{t}$ through gradient-descent type updates. We have formalised the resulting training procedure in Algorithm 1, and, given the close relationship between Theorem 2.1 and the Bismuth-Elworthy-Li (BEL) formula from Malliavin calculus (see Section 2.2), we refer to these methods as BEL-algorithms. Importantly, the score process 8 can be simulated efficiently by directly updating $J_{t\mid s}^{\top}(\sigma_{t}(X_{t})^{\top})^{-1}\,\mathrm{d}B_{t}$ along 7, without simulating the full dynamics of $J_{t\mid s}$ (see Appendix A for details). The main hyperparameter in Algorithm 1 is the matrix-valued function $\alpha:[0,T]\rightarrow\mathbb{R}^{n\times n}$; based on variance and stability considerations we give guidance on its choice in Section 3. Furthermore, specific choices of $\alpha$ allow us to connect existing methods to the framework of Theorem 2.1; those are explained in Sections 3.2.1 and 3.2.2.
In the remainder of this section we prove Theorem 2.1 and illustrate its connection to integration by parts, as hinted at in 5. Implementation details are deferred to Section 4.
Algorithm 1 BEL - Training Step
0:  $\alpha:[0,1]\to\mathbb{R}^{n\times n}$, initial condition $x_{0}$, batch size $N$, current drift approximation $u^{\theta}$, time grid $\{t_{0},t_{1},\ldots t_{M}\}$. Initialize
1:  for $i=1$ to $N$ do
2:     Sample a sample path $X$ with corresponding Brownian motion path $B$ from the SDE 6.
3:     Sample an observation $Y=G(X_{T})$ from 2.
4:     Compute the Monte Carlo estimator $\mathcal{S}_{s}$ using 8 (for details see Algorithm 2) along the path $(X,B)$.
5:     Calculate the single-path loss
$l_{i}(\theta)=\sum_{j=1}^{M-1}\|u_{t_{j}}^{\theta}(X_{t_{j}};Y)-\mathcal{S}_{t_{j}}(X,B)\|^{2},$
6:  end for
7:  Summ for the full-batch loss
$\mathcal{L}^{M}(\theta)=\sum_{i=1}^{N}l_{i}(\theta).$
8:  Take a gradient step on $\mathcal{L}^{B}(\theta)$ with your favourite optimiser.
2.1 Doob’s $h$-transform and Conditional Scores
As a first step towards Theorem 2.1, we recall the fact that the diffusion bridge drift in 10 can be expressed in terms of conditional scores (Rogers & Williams, 2000, p. 83):
Proposition 2.2 (Doob’s $h$-transform).
For $y\in\mathbb{R}^{d}$, let
$u_{t}^{*}(x;y)\coloneqq\sigma_{t}(x)\sigma_{t}(x)^{\top}\nabla_{x}\log p_{T\mid t}(Y=y\mid X_{t}=x),$ (11)
where $p(Y=y\mid X_{t}=x)$ is the probability of $Y=y$ given $X_{t}=x$. Then the corresponding controlled diffusion (10) reproduces the conditional law of (6), given $Y=y$.
Similar closed-form formulations are available in the context of optimal control (Nüsken & Richter, 2021, Section 2.2) and time reversals (Boffi & Vanden-Eijnden, 2024; Song et al., 2021; Ho et al., 2020). Note that for $t=T$ $\nabla_{x}\log p_{T|t}(Y=y|X_{T}=x)=\nabla_{x}\log g(x;y)$. As a consequence, if $g$ is singular, (11) becomes singular (and possibly numerically unstable) as $t\rightarrow T$. We deal with this issue in the next subsection.
2.2 Malliavin Calculus and Integration by Parts
While Proposition 2.2 in principle identifies the desired control vector field, the right-hand side of (11) will be unavailable in all but simple toy examples. The following result provides a formula that is amenable to Monte Carlo simulation:
Proposition 2.3 (Generalised Tweedie formula).
The conditional score is given by the conditional expectation of the score process,
$\begin{split}\nabla_{x}\log p_{T\mid t}&(Y=y\mid X_{t}=x)\\ &=\mathbb{E}\left[\mathcal{S}_{t}\mid X_{t}=x,Y=y\right]\end{split}$ (12)
for all $t\in[0,T)\,\text{and }x\in\mathbb{R}^{n},y\in\mathbb{R}^{d}.$
Proof sketch.
See Appendix D.1 for full details. We rely on two main ideas:
Firstly, to express the transition probability in terms of an expectation, we use the fact that whenever a random variable $X$ admits a smooth Lebesgue density $p_{X}$, we have the representation
$p_{X}(x)=\mathbb{E}[\delta_{x}(X)],$ (13)
see Duistermaat et al. (2010) for a rigorous general account or Watanabe (1987, Section 2.1) for the statement in the context of Malliavin calculus.
Secondly, as outlined in the introduction, we elevate the integration-by-parts formula (5) to Wiener space, i.e., to the space of sample paths of Brownian motion $(B_{t})_{0\leq t\leq T}$. To this end, we introduce the Malliavin derivative $D_{t}$, which represents differentiation with respect to the infinitesimal noise increment $\mathrm{d}B_{t}$: for a functional $F$ that depends on the realisation of $(B_{t})_{0\leq t\leq T}$, the Malliavin derivative is given by
$D_{t}F=\frac{\partial F}{\partial\,\mathrm{d}B_{t}}.$ (14)
A rigorous treatment of (14) requires the framework of calculus on Wiener space; see Nualart (2006). With (14) in place, the analogue of (5) on Wiener space becomes
$\mathbb{E}\left[\int_{0}^{T}D_{t}F\cdot g_{t}\,\mathrm{d}t\right]=\mathbb{E}\left[F\int_{0}^{T}g_{t}\cdot\mathrm{d}B_{t}\right],$ (15)
for appropriate choices of $F$ and the stochastic process $g_{t}$.
We will see in Proposition D.2 that we can express $\nabla\log p(Y=y|X_{t}=x)$ as a conditional expectation of $\nabla\log p(X_{T}=x_{T}|X_{t}=x)$. This simplifies the problem to inferring the diffusion bridge term $\nabla\log p(X_{T}=x_{T}|X_{t}=x)$, which we do now.
Using (13), we write
$\nabla_{x}\log p_{T\mid t}(X_{T}\!=\!x_{T}\mid X_{t}\!=\!x)\!=\!\frac{\nabla_{x}\mathbb{E}[\delta_{x_{T}}(X_{T})\mid X_{t}=x]}{p_{T\mid t}(X_{T}\!=\!x_{T}\mid X_{t}\!=\!x)},$
and, using the chain rule, we further obtain
$\nabla_{x}$ $\mathbb{E}[\delta_{x_{T}}(X_{T})\mid X_{t}=x]$
$=-\mathbb{E}[\nabla_{x_{T}}\delta_{x_{T}}(X_{T})J^{\top}_{T\mid t}\mid X_{t}=x].$
Here, $\nabla_{X_{t}}X_{T}=J_{T\mid t}$, as explained in Section 2, and the gradient $\nabla_{x_{T}}\delta_{x_{T}}$ is understood in the sense of distributions (Duistermaat et al., 2010, Chapter 4). To complete the proof, it remains to eliminate the derivative on $\delta_{x_{T}}$ via the integration by parts formula (15)—the left-hand side of (15) precisely explains the appearance of the stochastic integral in the score process (8). For this, we need to convert the conventional gradient $\nabla_{x_{T}}$ into the Malliavin derivative $D_{t}$. This is achieved using the matrix-valued function $\alpha$, as detailed in Appendix D.1. In fact, there are infinitely many ways to carry out this conversion, each determined by a particular choice of $\alpha$. We note that the proof strategy, particularly the conversion of $\nabla_{x_{T}}$ into $D_{t}$ and the use of integration by parts, closely resembles the proof of the Bismuth-Elworthy-Li (BEL) formula from Malliavin calculus (Bismut, 1984; Elworthy & Li, 1994), which inspired the name of Algorithm 1. ∎
Remark 2.4.
Note that $\nabla\log p_{t}$ contains the Lebesgue-density $p_{t}$ and therefore does not naturally carry over to infinite dimensions or manifolds. However, the expression (8) as well as (12) can also be made sense of in infinite dimensions or for diffusions on manifolds. Therefore, these are natural expressions to use for conditioning infinite as well as manifold-valued diffusions (Baker et al., 2024).
Remark 2.5.
For linear SDEs Proposition 2.3 can be simplfied to the celebrated Tweedie formula (Efron, 2011), given, e.g., by
$\nabla_{x_{T}}\log p_{T}(X_{T}=x_{T})=\tfrac{1}{T}(\mathbb{E}[X_{t}\mid X_{T}=x_{T}]-x_{T}).$
Similarly, the denoising score matching objective for diffusion models (Song et al., 2020) can be seen as an instance of 9 for a specific choice of $\alpha$. This also allows us to derive novel regression targets for diffusion models. Lastly, similar to Remark 2.5, the obtained formulas carry over to the infinite dimensional as well as manifold-valued setting and are natural targets to generalize diffusion models to these settings (Pidstrigach et al., 2024; De Bortoli et al., 2022). For details on this, see Appendix B.
2.3 Amortised Bridges and Forward-KL
While Propositions 2.2 and 2.3 together provide an expression for the desired control vector field $u_{t}^{*}$, the left-hand side of (12) requires access to the targeted distribution of (6), conditioned on $X_{t}=x$ and $Y=y$; the construction might thus appear circular. Fortunately, taking expectations over $X_{t}$ and $Y$ in (12) not only recovers the straightforward-to-simulate reference diffusion (6) as a mixture of its bridges, but also allows us to amortise the learning procedure, inferring all the bridges simultaneously.
The following proof of our main result reflects this idea:
Proof of Theorem 2.1.
Since $u$ is a function of $X_{t}$ and $Y$, and since the conditional expectation is the minimizer of the $L^{2}$ distance (Klenke, 2013, Corollary 8.17), the minimizer of
$\mathcal{L}^{t,y}_{\text{local}}(u):=\mathbb{E}\left[\left\|u_{t}(X_{t},Y)-\mathcal{S}_{t}\right\|^{2}\right],$ (16)
is given by the conditional expectation of $\mathcal{S}_{t}$ with respect to $X_{t}$ and $Y$, i.e.
arg $\text{arg min}~{}\mathcal{L}^{t}_{\text{local}}(u)=\mathbb{E}[\mathcal{S}_{t}|X_{t},Y]=\nabla_{X_{t}}\log p(Y|X_{t}).$ (17)
Here, the second equality holds because of Proposition 2.3. Finally, by Proposition 2.2, $\nabla_{X_{t}}\log p(Y|X_{t})$ is indeed equal to the control term. By exchanging the expectation with the integral in 9, we get that $u$ is equal to the optimal control $\nabla_{X_{t}}\log p(Y|X_{t})$ for each $t$.
Remark 2.6 (KL interpretation).
The time-integrated local loss can be interpreted in terms of the amortized $\mathrm{KL}$-divergence between measures on path space,
$\int_{0}^{T}\mathcal{L}^{t}_{\mathrm{local}}(u_{t})\,\mathrm{d}t=\mathbb{E}\left[\mathrm{KL}(\mathbb{P}^{y}\mid\mathbb{P}^{u})\right]+C$ (18)
with a constant $C>0$ that does not depend on $u$. Here, the expectation is taken with respect to $Y$ and $\mathbb{P}^{y}$ refers to the distribution of trajectories induced by the diffusion conditioned on $Y=y$. The measure $\mathbb{P}^{u}$ refers to the distribution associated to the current control $u$. The forward-$\mathrm{KL}$ divergence in (18) is mode-covering (Naesseth et al., 2020), which is a desirable property for conditional generation and rare event simulation. Furthermore, the $\mathrm{KL}$-representation (18) allows us to connect the BEL-framework from Algorithm 1 to the previous works by Heng et al. (2022) and Baker et al. (2025). However, the former requires additional numerical overhead in learning backward processes, whereas the latter relies on additional simulations for a Feynman-Kac type construction. The accuracy of these approaches is evaluated experimentally in Section 4. Finally, we remark that (18) distinguishes our approach from the work by Du et al. (2024), who use forward-$\mathrm{KL}$, together with subsequent (Gaussian) approximations. For background on the “measures on path space” perspective, see Nüsken & Richter (2021), and for a proof of 18 see Section D.2.
3 On the Choice of $\alpha$
Theorem 2.1 provides a whole class of loss functions for diffusion bridge simulation: each choice of $\alpha$ yields a slightly different approach to approximating the drift in (11), providing considerable flexibility. We now exploit this flexibility for two purposes: in Section 3.1, we derive an optimal choice of $\alpha$ under a simplified setting, and in Section 3.2, we connect our bridge losses to existing algorithms used in stochastic optimal control and diffusion bridge simulation.
3.1 Optimal $\alpha$ for Reduced Variance
Each choice of $\alpha$ in Theorem 2.1 implies a different regression target for the neural network. While all of these targets have the same mean, namely the diffusion bridge drift in (11), they differ in variance:
$\text{Var}(\mathcal{S}_{s})=\text{Var}\left(A^{-1}_{T\mid s}\int_{s}^{T}\alpha_{t}^{\prime}J_{t\mid s}^{\top}(\sigma_{t}(X_{t})^{\top})^{-1}\mathrm{d}B_{t}\right).$
Since the algorithm works by generating independent samples from $\mathcal{S}_{s}$ and then regressing against those, a lower variance of $\mathcal{S}_{s}$ is expected to lower the variance of the gradients. In the following result we compute and optimise the variance in a simplified setting, providing guidance for the choice of $\alpha$:
Lemma 3.1.
Set $Y=X_{T}$, $T=1$, $n=1$, $b=0$ and $\sigma=1$, i.e., $X_{t}$ is a one-dimensional Brownian motion conditioned on $X_{0}=x_{0}$ and $X_{1}=x_{1}$. Then the variance of the Monte-Carlo estimator
$\nabla\log p_{1\mid 0}(B_{1}=x\mid B_{0}=x_{0})\approx\int_{0}^{1}\alpha_{t}^{\prime}J^{\top}_{t\mid 0}(\sigma_{t}^{-1})^{\top}\mathrm{d}B_{t}$
letting $d=x_{1}-x_{0}$, is given by
$\frac{1}{\alpha_{1}-\alpha_{0}}\!\left(\int_{0}^{1}\left({\alpha^{\prime}_{t}}\right)^{2}\mathrm{d}t+\!\int_{0}^{1}\frac{(\alpha_{1}-\alpha_{t})^{2}}{(1-t)^{2}}\mathrm{d}t+d^{2}\!\right)\!,$ (19)
assuming that $\frac{\alpha_{1}-\alpha_{s}}{\sqrt{1-s}}\rightarrow 0$ as $s\rightarrow 1$. The variance is minimised for the choice
$\alpha^{\prime}_{t}=1-(1-t)^{\frac{1}{2}\left(1+\sqrt{5}\right)}.$ (20)
The proof of Lemma 3.1 can be found in Appendix E, and Figure 2 shows the derivative $\alpha_{t}^{\prime}$ of the optimal variance-reducing choice in (20). Interpreted as weights in the score process (8), we see that (20) weights the initial increments of the Brownian motion relatively highly in comparison to increments closer to the terminal time. The choice of $\alpha$ will further be explored in Section 4.
Figure 2: The variance-optimal weighting $\alpha_{s}^{\prime}$ (see Lemma 3.1) for the simplified setting in which $X_{t}$ is a conditioned Brownian motion.
3.2 Connections to other Algorithms
3.2.1 Reparametrisation
Inspired by the reparametrisation trick in variational inference (Kingma et al., 2019, Section 2.4), Domingo-Enrich et al. (2024b) derived a novel methodology for stochastic optimal control. A variant of it–adapted to our setting–can be recovered from Theorem 2.1 by a specific choice of $\alpha$:
Lemma 3.2.
Let $M:[0,T]\rightarrow\mathbb{R}^{n\times n}$ be a matrix-valued differentiable function such that $M_{0}=\operatorname{Id}$ and $M_{T}=0$. Then the score process from (8) has the equivalent representation
$\mathcal{S}_{s}=\int_{s}^{T}(M_{t}\nabla b_{t}^{\top}(X_{t})-M_{t}^{\prime})(\sigma_{t}(X_{t})^{\top})^{-1}\,\mathrm{d}B_{t}.$ (21)
Proof.
The correspondence between (8) is via the choice $\alpha_{t}=J_{t\mid s}^{-1}M_{t}$; see Appendix D.3. ∎
The significance of Lemma 3.2 is that (i) it extends the reparameterisation trick to the singular reward setting and (ii) it gives conceptual insights into the choices of $\alpha_{t}$ and $M_{t}$ in the respective methods. For further intuition into the role of $M_{t}$ we refer to Domingo-Enrich et al. (2024b) and Appendix D.3.
3.2.2 Gaussian Approximations
Another method to learn $\nabla\log p(Y=y\mid X_{s}=x_{s})$, identified as the conditional drift in Proposition 2.3, is to approximate the transition densities as Gaussian and regress against their score. We now derive methods based on a Gaussian approximation and show how they can be reformulated as BEL algorithms for specific choices of $\alpha$. Moreover, the BEL algorithm provides deeper insight into the approximation error introduced by the Gaussian assumption.
Rather than regressing directly against the target, we first observe that one can regress against the optimal drift for a small timestep:
Lemma 3.3.
For any $t\geq s$,
$\mathcal{L}^{\text{GA}}(u,s,t)$ (22)
$~{}~{}~{}~{}~{}\coloneqq\mathbb{E}\big{[}\|u_{s}(X_{s},X_{T})-\nabla_{X_{s}}\log p(X_{t}\mid X_{s})\|^{2}\big{]}$
is given by the diffusion bridge $u^{*}_{t}(x_{t},x_{T})$, see (11).
Thus, optimizing $\mathcal{L}^{\text{GA}}$ for each $s$ yields the diffusion bridge drift. To achieve this, one must select a $t\geq s$ for each $s$. After discretizing the time domain into $\{t_{0},t_{1}=t_{0}+\delta t,\ldots,t_{N}=t_{0}+N\delta t\}$, a suitable loss function is:
$\mathcal{L}^{\text{GA}}(u)=\sum_{i=1}^{N-1}\mathcal{L}^{\text{GA}}(u,t_{i},t_{i}+\delta t).$ (23)
For small $\delta t$, the transition density can be approximated by a Gaussian, for example, via an Euler-Maruyama step:
$p(X_{t+\delta t}\mid X_{t})$ $\approx\mathcal{N}(X_{t}+\delta t~{}b_{t}(X_{t}),\delta t~{}a_{t}(X_{t})),$ (24a)
$a_{t}(X_{t})$ $=\sigma_{t}(X_{t})\sigma_{t}(X_{t})^{T}$ (24b)
This provides an explicit expression for $\nabla_{X_{t}}\log p(X_{t+\delta t}\mid X_{t})$, which can be used for regression. In Heng et al. (2022), the authors applied a similar Gaussian approximation to the time-reversal of a diffusion to learn bridges of the time-reversed process.
However, this approximation is performed in the density domain $p(X_{t+\delta t}\mid X_{t})$, and its impact on the accuracy of $\nabla\log p(X_{t+\delta t}\mid X_{t})$ is unclear. By the generalized Tweedie formula (see Proposition 2.3), we obtain the explicit relation:
$\nabla\log p_{t}(X_{t+\delta t}\mid X_{t})=\mathbb{E}[\mathcal{S}_{s}\mid X_{t},X_{t+\delta t}],$ (25)
where $\mathcal{S}_{s}$ is defined in (8) with terminal time $T=t+\delta t$.The Gaussian approximation can now directly be related to a discretization scheme for $\mathcal{S}_{s}$ in (25). We prove this in the case of $\sigma=1$:
Lemma 3.4.
Approximating $p(X_{t+\delta t}\mid X_{t})$ by a Gaussian (24a) and regressing against its score is equivalent to a choosing $\alpha^{\prime}_{s}=1_{[t,t+\delta t]}$ and approximating the stochastic integral in $\mathcal{S}_{s}$ (8) as
$\int_{t}^{t+\delta t}J_{s\mid t}\mathrm{d}B_{s}\approx J_{t+\delta t\mid t}(B_{t+\delta t}-B_{t}),$ (26)
and furthermore approximating $J_{t+\delta t\mid t}$ by an Euler-Maryuama step on (7):
$J_{t+\delta t\mid t}\approx\operatorname{Id}+\delta t\nabla b(t,X_{t}).$ (27)
The proofs for this section can be found in Section D.4.‡
3.3 Summary of Choices
Based on the above discussion we will now summarise some choices for the function $\alpha_{t}$ in Theorem 2.1. Each choice leads to a different loss function, and we compare the different choices empirically in Section 4.
BEL optimal: The first choice is setting $\alpha$ as in LABEL:{lemma:variance-bel-bm}. Although this only gives optimal variance in the specific case of Brownian motion, it still may do well in other problem settings, especially those similar to Brownian motions.
BEL first: Based on Lemma 3.3 we can set $\alpha_{s}$ such that $\alpha^{\prime}_{s}={1}_{[s,s+\Delta s]}$. In this case, $\alpha_{T}-\alpha_{s}=\Delta s$. This means we are only using local information to approximate the score.
BEL average: Another choice is setting $\alpha_{s}=s$. This leads to the traditional Bismut-Elworth-Li formula (Bismut, 1984; Elworthy & Li, 1994).
BEL last: This algorithm uses $\alpha^{\prime}_{s}={1}_{[T-\Delta t,T]}$. This is similar to the stochastic optimal control setting, where the gradient of the target function is propagated back through the whole trajectory.
Reparametrisation: Corresponding to the discussion on the reparametrisation trick in Lemma 3.2, we set $\alpha_{t}=J_{t\mid s}M_{t}$, with $M_{t}=\frac{T-t}{T}\operatorname{Id}$.
Figure 3: In (a), we plot the double-well potential 31 for $v=5$. In (b), we sample 1,000 paths from the unconditioned SDE (30) and observe that they remain confined to a single potential minimum, failing to transition between them. This highlights the inherent difficulty of the problem.
4 Experiments
In this section, we do an empirical evaluation of our methods. Full experimental details are provided in Appendix C. In Section 4.1 and Section 4.2 we study the case of diffusion bridges, i.e. $Y=X_{T}$, or a Dirac-delta reward function $g$, since this can be seen as the most challenging setup: In Section 4.1, we conduct experiments where the true transition densities are available for evaluation. In Section 4.2, we demonstrate our method on bridges of stochastic shape processes, which have applications in biology. We also compare to related methods (Heng et al., 2022) and (Baker et al., 2025) and show favourable results. In Section 4.3 we apply our methodology to diffusion models or equivalently, flow matching algorithms.
4.1 Controlled Environment Experiments
In this subsection, we introduce two experiments where the true diffusion bridge drift can be computed, allowing us to evaluate our methods against the ground truth. We explain the experiment setup in Section C.1.1.
4.1.1 Methodology
For the experiments in Section 4.1.2 and Section 4.1.3, we consider one-dimensional SDEs, simulated independently across all dimensions. Trajectories are conditioned to start at $y_{\text{init}}=(1,1,\dots,1)\in\mathbb{R}^{D}$ and end at $y_{\text{final}}=(-1,1,\dots,1)\in\mathbb{R}^{D}$, modifying the first coordinate.
This setup isolates the effect of the rare event from the dimensionality of the state space. Conditioning all coordinates would cause the probability of the transition event—being the product of independent one-dimensional transition probabilities—to scale as $p^{D}$, where $p$ is the probability of the event in one dimension. In contrast, by constraining only the first coordinate, we maintain an approximately constant event probability whilst varying the problem dimensionality.
4.1.2 Brownian Motion
The first experiment we consider involves conditioning a Brownian motion:
$\mathrm{d}X_{t}=\mathrm{d}B_{t}.$ (28)
For this SDE, the diffusion bridge drift (11) has a closed-form solution given by:
$\nabla\log p(X_{1}=x\mid X_{t}=x_{t})=-\frac{x_{t}-x}{1-t}.$ (29)
This allows us to compare our methods against the true bridges. The results are presented in Figure 8. We observe that the reparametrisation trick algorithm performs well for the one-dimensional Brownian motion, while the BEL average achieves the best performance in $10$ dimensions. For an explanation of the metrics see Section C.1.2
Figure 4: Panel (a) shows ground truth paths from 30, conditioned on transitioning from state $1$ to state $-1$. Panels (b), (c) and (d) present the paths generated by the BEL-first, BEL-last and BEL-optimal respectively.
4.1.3 Double-Well
In this experiment, we consider the double-well problem as described by Nüsken & Richter (2021). This model is given by the SDE:
$\mathrm{d}X_{t}$ $=-\nabla U_{v}(X_{t})\mathrm{d}t+\mathrm{d}B_{t},$ (30)
$U_{v}(x)$ $=v(x^{2}-1)^{2}.$ (31)
For $v=5$, we visualize the potential in Figure 3. The potential exhibits two minima at $x=-1$ and $x=1$, which correspond to metastable states. Since the drift term in 30 drives trajectories toward the minima of $U_{v}$, transitions between $x=-1$ and $x=1$ are rare events, with their probability decreasing exponentially as the potential barrier height $v$ increases (Kramers, 1940; Berglund, 2013). We illustrate this by plotting $1000$ sample paths of the unconditioned process in Figure 3, none of which crosses the barrier.
To obtain a ground truth for this example, we numerically estimate $f(x)=p(X_{1}=-1\mid X_{s}=x)$ and compute its logarithmic gradient. We then compare the paths of the process under the true drift with our estimates in Figure 4.
Additionally, we extend the experiment to higher-dimensional settings as described in Section 4.1.1. The first and second marginals of a $10$ dimensional process are shown in Figure 5.
Finally, we quantitatively compare the performance of different algorithms in Table 4. We observe that BEL Last and the Reparameterisation Trick exhibit the lowest performance, aligning with the empirical experience of the authors. For an explanation of the metrics, see Section C.1.2. In Figure 7 we also provide plots for the reparametrisation trick.
Figure 5: The first and second coordinates of a 10-dimensional process sampled from the SDE in 30 are shown in (a). The process is conditioned on transitioning from $1$ to $-1$ in the first coordinate while remaining at $1$ in all others. Panels (b), (c), and (d) depict the paths generated by the BEL-first, BEL-average, and BEL-optimal algorithms, respectively. All plots include underlying contour lines representing the level sets of the potential.
4.2 Shape Processes
We demonstrate our methodology on stochastic shape processes (Arnaudon et al., 2023), which are used in computational anatomy to model morphological changes in human organs due to disease (Arnaudon et al., 2017). In evolutionary biology, they help analyse morphometric changes in species, such as how butterfly wing shapes evolve along phylogenetic trees (Baker et al., 2024).
Following the setup in (Sommer et al., 2021), we consider a shape represented by $x_{0}\in\mathbb{R}^{2N}$, where $N$ points discretise a two-dimensional shape. Let $\{y^{j}\}_{j=1}^{M}\subset\mathbb{R}^{2}$ be equidistant points, and $\{B^{j}\}_{j=1}^{M}$ be independent Brownian motions in $\mathbb{R}^{2}$. The SDE governing the shape evolution is given by
$\mathrm{d}x^{i}_{t}$ $=\sum_{j=1}^{M}k(y^{j},x^{i}_{t}),\mathrm{d}B^{j}_{t},\quad 1\leq i\leq N,$ (32)
$k(x,y)$ $=\kappa\frac{\|x-y\|_{2}^{2}}{\beta},$ (33)
where $k$ is a Gaussian kernel with parameters $\kappa,\beta\in\mathbb{R}$. For each $t$, the map $x_{0}\mapsto x_{t}$ is a diffeomorphism, ensuring that nearby points remain highly correlated. This property makes learning bridges particularly challenging. An example trajectory of the unconditioned process is shown in Figure 9.
We apply our method, BEL average (i.e. $\alpha_{t}=t$), to learn bridges of the shape SDE, conditioning the process on a circle of radius 1.5. We compare our approach to the methods of Heng et al. (2022) and Baker et al. (2025). The latter uses “adjoint” processes derived from a Feynman-Kac representation for the conditional expectation in 11, while Heng et al. (2022) estimate the time-reversed bridge process by regressing against the time-reversed diffusion bridge drift for small time steps, following Lemma 3.3. However, due to the time-reversal their drift term involves the gradient $\nabla_{x_{t}}p_{t\mid s}(X_{t}=x_{t}\mid X_{s}=x_{s})$, where differentiation is taken with respect to the final time point rather than the initial one.
Our findings are given in Table 1 and show that BEL average outperforms competing methods. The next best method is Heng et al. (2022), which can be interpreted as first applying BEL to the time-reversed process, approximating the transition densities via Gaussians (24a), and using a fixed start point instead of amortisation. However, since their approach involves time-reversal, they do not take gradients of the drift or diffusion terms. BEL average not only uses this gradient information, but non-local information about the score.
Method Dist
BEL average $0.085$
Time reversal $0.090$
Adjoint paths $0.498$
Untrained $1.396$
Table 1: A comparison between different methods for learning bridges for shape processes. We see that our proposed method BEL average outperforms other existing methods.
4.3 Image Experiments
We train a diffusion model using a flow matching loss with a U-Net architecture on the Fashion-MNIST dataset (Xiao et al., 2017).
Remark 4.1.
The model is trained deterministically using the flow matching loss. We then apply the memoryless schedule from Domingo-Enrich et al. (2024a) to reinterpret the trained model as an equivalent SDE (or diffusion model).
Next, we condition the resulting SDE to produce images with a specified upper-left corner. Importantly, this conditioning does not require adding artificial noise to the observation. More specifically, the law of $X_{T}$ is by design the data distribution. Then we condition on $Y=G(X_{T})$ where G selects the upper left corner of the image. Once trained we can therefore choose an arbitrary upper left corner $y$ and sample from images matching this, using the learned additional drift $u(X_{t},y)$. Note we only need train once and then can sample using any corner. We demonstrate the result by conditioning on the upper-left corner of a jacket image in Figure 6.
Remark 4.2.
When both the forward and reverse dynamics of an SDE are known, one can construct bridges, as shown in Heng et al. (2022). Diffusion models form a special case of this setting: not only are both directions available, but the reverse dynamics are particularly easy to simulate. This property has been leveraged in Denker et al. (2024) to simplify the methodology of Heng et al. (2022). However, both approaches implicitly assume that the learned score is exact. In contrast, our method makes no such assumption. Since we do not require access to the reverse dynamics, we directly learn the correct conditioning for the learned score—even when it is inexact.
Figure 6: In panel (a), the top-left image shows the ground truth used for conditioning. The remaining 15 images are samples generated by the diffusion model conditioned on the upper-left quarter of the ground truth image. Panel (b) displays only the conditioning inputs: again, the top-left image is the ground truth, while the others show the corresponding conditioned quarters used for generation.
5 Conclusion
In this work, we have introduced a novel class of loss functions for estimating the drift of a conditioned diffusion process. Our approach leverages Malliavin calculus and integration by parts to handle singular losses, enabling a more stable and accurate estimation. This framework not only leads to a generalized Tweedie’s formula but also unifies several existing methods under a common theoretical perspective. Empirically, we demonstrated that our loss functions outperform competing algorithms in the simulation of shape processes, and scales to high dimensions on image experiments. Future work includes extending our methodology to a broader range of applications and further investigating the role of $\alpha$, either through analytical techniques or by optimising it.
Table 2: Dimension 1
Loss MV Dist
BEL average $8.1\times 10^{-1}$ $1.6\times 10^{-1}$
BEL first $2.9\times 10^{-1}$ $8.4\times 10^{-2}$
BEL last $1.1\times 10^{0}$ $1.1\times 10^{0}$
BEL optimal $8.3\times 10^{-1}$ $1.9\times 10^{-1}$
Reparametrization Trick $9.2\times 10^{-1}$ $2.1\times 10^{-1}$
Table 3: Dimension 10
Loss MV Dist
BEL average $13.4\times 10^{-}$ 1 $13.0\times 10^{-}$ 1
BEL first $12.1\times 10^{-}$ 1 $13.3\times 10^{-}$ 1
BEL last $15.9\times 10^{-}$ 1 $1.1\times 10^{0}$
BEL optimal $13.1\times 10^{-}$ 1 $13.0\times 10^{-}$ 1
Reparametrization Trick $15.5\times 10^{-}$ 1 $15.2\times 10^{-}$ 1
Table 4: Performance of our proposed algorithms for conditioning the double well SDE (see Section 4.1.3). The best-performing algorithm metrics are marked in red.
Acknowledgments
This work is partially supported by the Novo Nordisk Foundation grant NNF18OC0052000, a research grant (VIL40582) from VILLUM FONDEN, and UCPH Data+ Strategy 2023 funds for interdisciplinary research. GD and JP were supported by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1].
Impact Statement
This work has connections to diffusion models, and could be used in that setting. As such, it shares the societal and ethical implications of generative modelling.
Appendix A Adjoint SDE for Calculating $\mathcal{S}_{s}$
Algorithm 2 Adjoint SDE Method for Calculating $\mathcal{S}_{s}$
0:  Simulation path $\{X_{t}\}$, Brownian increments $\{\delta B_{t}\}$, where $t\in\{0,\delta t,\ldots,T\}$.
1:  Initialise $\tilde{\mathcal{S}}_{T}=0$.
2:  for $t=T-\delta t$ to 00 (backwards in steps of $\delta t$do
3:     Calculate $\mathcal{D}_{t}$ via 36.
4:     Update $\tilde{\mathcal{S}}_{t}=\tilde{\mathcal{S}}_{t+\delta t}+\mathcal{D}_{t}$.
5:  end for
6:  Compute score process $\mathcal{S}_{s}$ for $s\in\{0,\delta t,\ldots,T\}$ via 37.
We employ an adjoint stochastic differential equation (SDE) method, inspired by adjoint ordinary differential equations (ODEs) (Kidger, 2022; Pontryagin, 2018; Chen et al., 2018), to compute $\mathcal{S}_{s}$. Specifically, we define the auxiliary variable $\tilde{\mathcal{S}}_{s}$ as:
$\tilde{\mathcal{S}}_{s}:=\int_{s}^{T}J_{t\mid s}^{\top}(\sigma_{t}(X_{t})^{\top})^{-1}\,\mathrm{d}B_{t},$ (34)
which corresponds to the score process in equation (8) when $\alpha^{\prime}\equiv 1$, disregarding the normalisation factor $A$.
We assume access to the following data obtained from a discretised Euler-Maruyama simulation of the SDE (6):
  • $X_{0},X_{\delta t},\ldots,X_{T}$: The states of the discretised SDE.
  • $\delta B_{0},\delta B_{\delta t},\ldots,\delta B_{T-\delta t}$: The Brownian increments used in the simulation, where $X_{t+\delta t}=X_{t}+\delta tb_{t}(X_{t})+\sigma(X_{t})\delta B_{t}$.
Note that $\tilde{\mathcal{S}}_{T}=0$. We can recursively compute $\tilde{\mathcal{S}}_{s}$ from $\tilde{\mathcal{S}}_{s+\delta t}$ using:
$\tilde{\mathcal{S}}_{s}=\int_{s}^{T}J_{t\mid s}^{\top}\left(\sigma_{t}^{\top}\right)^{-1}(X_{t})\,\mathrm{d}B_{t}=J_{s+\delta t\mid s}^{\top}\tilde{\mathcal{S}}_{s+\delta t}+\int_{s}^{s+\delta t}J_{t\mid s}^{\top}\left(\sigma_{t}^{\top}\right)^{-1}(X_{t})\,\mathrm{d}B_{t},$
making use of the semigroup property $J_{t|s}=J_{t|s+\delta t}J_{s+\delta t|s}$.
Approximating the integral term, we obtain
$\tilde{\mathcal{S}}_{s}\approx J_{s+\delta t\mid s}^{\top}\tilde{\mathcal{S}}_{s+\delta t}+J_{\tau\mid s}^{\top}\left(\sigma_{t}^{\top}\right)^{-1}(X_{s})\delta B_{s},$ (35)
where $\tau$ can be any number in $\tau\in[s,s+\delta t]$. For $\tau=s$, we have $J_{s\mid s}=\text{Id}$. The term $J_{s+\delta t|s}$ can be approximated using an Euler-Maryuama step on 7, leading to $J_{s+\delta t\mid s}\approx\text{Id}+\delta t\nabla b_{s}(X_{s})+\nabla\sigma_{s}(X_{s})\delta B_{s}$.
Remark A.1.
Here, the term $\nabla\sigma_{t}(X_{t})~{}\delta B_{t}$ is meant as the derivative with respect to $x$ of the map
$x\in\mathbb{R}^{n}\mapsto\sigma_{t}(x)\delta B_{t}\in\mathbb{R}^{n}.$
This allows us to compute the difference term:
$\mathcal{D}_{s}:=\tilde{\mathcal{S}}_{s}-\tilde{\mathcal{S}}_{s+\delta t}=\int_{s}^{s+\delta t}J_{t\mid s}^{\top}(\sigma_{t}(X_{t})^{\top})^{-1}\,\mathrm{d}B_{t}\approx(\delta t\nabla b(X_{s})^{\top}+\nabla\sigma(X_{s})^{\top}\delta B_{s})\tilde{\mathcal{S}}_{s+\delta t}+(\sigma_{t}^{\top})^{-1}(X_{s})\delta B_{s}.$ (36)
Crucially, note that these Jacobian-vector products are efficiently computed using reverse-mode autodifferentiation, avoiding explicit Jacobian computation.
The algorithm proceeds by initialising $\tilde{\mathcal{S}}_{T}=0$ and iteratively computing $\tilde{\mathcal{S}}_{t}$ backward in time for $t\in\{T-\delta t,\ldots,\delta t,0\}$. Finally, $\mathcal{S}_{s}$ is obtained as
$\mathcal{S}_{s}\approx A_{T\mid s}^{-1}\sum_{t=s,s+\delta t,\cdots}^{T-\delta t}\alpha_{t}^{\prime}\mathcal{D}_{t},$ (37)
restoring the normalisation that was dropped in (34).
Appendix B Denoising Score Matching and Tweedies Formula
Here we show how our results can be applied to derive a formula for the score $\nabla\log p_{t}(x)$ of a diffusion process. This could be done by applying Proposition 2.3 to a trivial conditioning $Y=0$ and treating the reverse dynamics of the SDE (6). However, it is instructive to quickly rederive the analogous equation in this simplified setting, which we do in this section. We will see that it generalises Tweedies formula.
Assume we have an SDE
$\mathrm{d}X_{t}=b_{t}(X_{t})\,\mathrm{d}t+\sigma_{t}(X_{t})\,\mathrm{d}B_{t},$ (38)
and denote by $p_{t}$ the density of $X_{t}$. Then we have the following representation:
Lemma B.1.
The score can be represented as
$\nabla\log p_{t}(x)=\mathbb{E}\left[\int_{0}^{t}(\sigma(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}|X_{t}=x\right].$
Here $\alpha_{s}^{\prime}$ is a function which satisfies $\int_{0}^{t}\alpha_{s}^{\prime}\,\mathrm{d}s=1$, and $J_{t|s}$ is the Jacobian which follows the flow (7).
In particular, for any such $\alpha$, the loss
$\mathcal{L}(u)=\mathbb{E}\left[\|u(X_{t})-\int_{0}^{t}(\sigma(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\|^{2}\right]$ (39)
has a unique minimiser given by $u(x)=\nabla\log p_{t}(x)$.
Proof.
By the chain rule for Malliavin derivatives we can write
$D_{s}\varphi(X_{t})=\nabla\varphi(X_{t})J_{t|s}\sigma_{s}(X_{s}),$
where
$J_{t|s}=\nabla_{X_{s}}X_{t}$
is the Jacobian. This implies that
$\nabla\varphi(X_{t})=D_{s}\varphi(X_{t})\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1},$
assuming that $\sigma_{s}$ is invertible. As this relationship holds for all $s$ we can integrate over $s$ to get
$\nabla\varphi(X_{t})=\int_{0}^{t}D_{s}\varphi(X_{t})\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1}\alpha_{s}^{\prime}\,\mathrm{d}s$
making use of the fact that $\alpha_{s}^{\prime}$ integrates to $1$. Taking expectations and using Malliavin integration by parts (see (15)), we arrive at
$\mathbb{E}[\nabla\varphi(X_{t})]=\mathbb{E}\left[\int_{0}^{t}D_{s}\varphi(X_{t})\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1}\alpha_{s}^{\prime}\,\mathrm{d}s\right]=\mathbb{E}\left[\varphi(X_{t})\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\right].$
Based on the framework from Watanabe (1987), we apply this to $\varphi=\delta_{x}$ the Dirac delta centred at $x\in\mathbb{R}^{n}$, and get
$\nabla\log p_{t}(x)$ $=\nabla_{x}\log\int p_{t}(z)\delta_{x}(z)\,\mathrm{d}z=\frac{1}{p_{t}(x)}\nabla_{x}\int p_{t}(z)\delta_{x}(z)\,\mathrm{d}z=\frac{1}{p_{t}(x)}\nabla_{x}\mathbb{E}[\delta_{x}(X_{t})]$
$=\frac{1}{p_{t}(x)}\mathbb{E}\left[\delta_{x}(X_{t})\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\right]=\mathbb{E}\left[\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}|X_{t}=x\right].$
Since the conditional expectation is the minimiser of the $L^{2}$ distance, (39) follows. ∎
Lemma B.2.
Assume that the forward SDE (38) is an Ornstein-Uhlenbeck process
$\mathrm{d}X_{t}=-\frac{1}{2}X_{t}\,\mathrm{d}t+\mathrm{d}B_{t}.$ (40)
Then for the choice $\alpha_{s}^{\prime}=e^{s-t}$ we get Tweedie’s formula:
$\nabla\log p_{t}(x)=\frac{1}{1-e^{-t}}\mathbb{E}\left[X_{t}-e^{-t/2}X_{0}|X_{t}=x\right],$ (41)
and, in particular, the loss (39) is simplifies to the well-known denoising score matching loss:
$\mathcal{L}(u)=\mathbb{E}\left[\left\|u(X_{t})-\frac{1}{1-e^{-t}}(X_{t}-e^{-t/2}X_{0})\right\|^{2}\right].$ (42)
Proof.
We have that
$J_{t|s}=e^{-(t-s)/2}\text{Id}.$
Therefore,
$\nabla\log p_{t}(x)=\mathbb{E}\left[\int_{0}^{t}(\sigma(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}|X_{t}=x\right]=\mathbb{E}\left[\int_{0}^{t}e^{\frac{t-s}{2}}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}|X_{t}=x\right].$
If we choose $\alpha_{s}=e^{s-t}$, then
$\nabla\log p_{t}(x)=\frac{1}{1-e^{-t}}\mathbb{E}\left[\int_{0}^{t}e^{-\frac{t-s}{2}}\mathrm{d}B_{s}|X_{t}=x\right].$ (43)
However, the solution of (40) is given by
$X_{t}=e^{-t/2}X_{0}+\int_{0}^{t}e^{-\frac{t-s}{2}}\,\mathrm{d}B_{t}.$
Plugging this into (43), we get
$\nabla\log p_{t}(x)=\frac{1}{1-e^{-t}}\mathbb{E}[X_{t}-e^{-t/2}X_{0}|X_{t}=x].$
Now, (42) follows again since the conditional expectation is the minimiser of the $L^{2}$ distance.
Appendix C Experimental Details
C.1 Controlled Environment Experiments
C.1.1 Experiment Setup
We discretised the time domain $[0,1]$ into $200$ equivariant grid points for the simulation and used an Euler-Maryuama scheme to simulate paths of the unconditioned SDE (6). We approximated $\mathcal{S}_{s}$ (8) by starting at the final increment of $\mathrm{d}B_{t}$ and then using an efficient adjoint method for SDEs to propagate the derivative information backwards. The full Jacobian matrix $J_{s\mid t}$ is never calculated.
For the algorithms BEL first and BEL last we used $\delta t=\frac{1}{200}$ (see Section 3.3).
We used a batch size of $2048$ and iterated through $20~{}000$ batches. We used the Adam optimizer and a neural network architecture which is loosely inspired by UNets. It projects the input data up to $256$ dimensions and then has fully connected layers of size $[256,128,64,32,64,128,256]$ with skip connections. The last layer is then a fully connected layer to the output dimension.
C.1.2 Metrics
Both metrics compare the simulated paths with paths from the ground truth when started in $y_{\text{init}}$ and conditioned on landing in $y_{\text{final}}$ (see Section 4.1.1), to see if the algorithms can approximate rare events even though they are trained on an amortised objective. We used $15~{}000$ simulations of the trained models to calculate the metrics.
Dist.
This metric calculates the average Euclidean distance of the final state of the path to the conditioned $y_{\text{final}}$.
MV.
This metric calculates the coordinate-wise mean and variance along the paths of the SDE. It then compares those to the coordinate-wise mean and variance of paths $m^{\text{gen}},v^{\text{gen}}$ simulated with the ground truth drift, and calculates
$\operatorname{MV}=\sqrt{\frac{1}{200}\sum_{i=1}^{200}\|m^{\text{gen}}_{i}-m_{i}\|^{2}+\|(v_{t}^{\text{gen}})^{1/2}-v_{t}^{1/2}\|^{2}},$ (44)
where $m_{i}$ and $v_{i}$ are the variance vectors at time $t=\frac{i}{200}$. The form of (44) is inspired by the Wasserstein-2 distance of two normal distributions.
C.1.3 Results
Here we provide the tables with the results for our controlled environment experiments for Brownian motion Figure 8.
We also provide figures for the reparametrisation trick for the double well experiments in Figure 7.
Figure 7: We plot the results from the double well experiments for the reparametrisation trick corresponding to (a) Figure 4, the one double well experiment and (b) Figure 5, the ten dimensional double well experiment.
Table 5: Dimension 1
Loss MV Dist
BEL average $7.6\times 10^{-2}$ $3.2\times 10^{-3}$
BEL first $8.6\times 10^{-2}$ $2.1\times 10^{-2}$
BEL last $1.5\times 10^{-1}$ $5.6\times 10^{-2}$
BEL optimal $1.1\times 10^{-1}$ $3.4\times 10^{-3}$
Reparametrization Trick $7.6\times 10^{-2}$ $2.2\times 10^{-3}$
Table 6: Dimension 10
Loss MV Dist
BEL average $14.5\times 10^{-}$ 1 $22.3\times 10^{-}$ 2
BEL first $14.5\times 10^{-}$ 1 $27.1\times 10^{-}$ 2
BEL last $14.8\times 10^{-}$ 1 $11.9\times 10^{-}$ 1
BEL optimal $14.5\times 10^{-}$ 1 $26.1\times 10^{-}$ 2
Reparametrization Trick $14.5\times 10^{-}$ 1 $29.4\times 10^{-}$ 2
Figure 8: Performance of various algorithms for conditioning a Brownian motion (see Section 4.1.2).
C.2 Shape Processes
For the kernel parameters in the SDE (32) we set $\kappa=0.1$ and $\beta=1.0$. For all methods we use the neural network and associated parameters in Yang et al. (2025) to train the model with the Adam optimiser. We discretise the time domain $[0,1]$ into $100$ equivariant grid points and use the Euler-Maruyama scheme to simulate paths of the SDEs. For each method, we train on a total of $102.400$ trajectories with a batch size of $128$. We compare to the time-reversal method (Heng et al., 2022) and the adjoint method (Baker et al., 2025) using the code provided by Baker et al. (2025).
To evaluate performance, we use the mean pointwise distance between the target shape $\{y_{i}\}_{i=1}^{N}$ and the final points of $M$ sampled bridges $\{x^{j}\}_{j=1}^{M}$:
$\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}\|x^{j}_{i}-y_{i}\|_{2}^{2}.$ (45)
We compute this metric over $M=512$ trajectories with $N=50$ points and report results in Table 1.
Figure 9: We plot one trajectory from the unconditioned stochastic shape process, started at a circle of radius 1.
Appendix D Proofs
D.1 Proof of Proposition 2.3
We start by proving the following preliminary result, which is a slight generalisation of the Bismuth-Elworthy-Li formula (Bismut, 1984; Elworthy & Li, 1994). Results in a similar spirit (although for transition densities instead of scores) can be found in Milstein et al. (2004).
Theorem D.1.
Let $\alpha:[0,T]\rightarrow\mathbb{R}^{n\times n}$ be a matrix-valued differentiable function such that $A_{T\mid s}\coloneqq\alpha_{T}-\alpha_{s}$ is invertible for all $s\in[0,T)$. For $T>s$ we have the representation formula
$\nabla_{x_{s}}\mathbb{E}[\varphi(X_{T})\mid X_{s}=x_{s}]=~{}\mathbb{E}\left[\varphi(X_{T})\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}\mid X_{s}=x_{s}\right]A_{T\mid s}^{-1}.$ (46)
Proof.
Case $Y=X_{T}$ (or $G=\text{Id}$). For any $s<T$ it holds that
$D_{s}\varphi(X_{T})=\nabla\varphi(X_{T})D_{s}X_{T}=\nabla\varphi(X_{T})J_{T\mid s}\sigma_{s}(X_{s})$
by the chain rule of Malliavin calculus (Nualart, 2006, Chapter 2). Therefore,
$\nabla\varphi(X_{T})=D_{s}\varphi(X_{T})\sigma_{s}(X_{s})^{-1}J_{T\mid s}^{-1},$ (47)
and
$\nabla_{x}\varphi(X_{T}^{x})=\nabla\varphi(X_{T})J_{T\mid 0}=D_{s}\varphi(X_{T})\sigma_{s}(X_{s})^{-1}J_{T\mid s}^{-1}J_{T\mid 0}=D_{s}\varphi(X_{T})\sigma_{s}(X_{s})^{-1}J_{s\mid 0}.$
Since (47) holds for any $s$, we can integrate along $\alpha_{s}^{\prime}$:
$\nabla_{x}\varphi(X^{x}_{T})=\int_{0}^{T}D_{s}\varphi(X_{T})\sigma_{s}(X_{s})^{-1}J_{s\mid 0}\alpha_{s}^{\prime}\mathrm{d}s(\alpha_{T}-\alpha_{0})^{-1}.$
We now apply the Malliavin integration by parts 15 to obtain
$\mathbb{E}[\nabla_{x}\varphi(X^{x}_{T})]$ $=\mathbb{E}\left[\int_{0}^{T}D_{s}\varphi(X_{T})\sigma_{s}(X_{s})^{-1}J_{s\mid 0}\alpha_{s}^{\prime}\mathrm{d}s(\alpha_{T})^{-1}\right]$
$=\mathbb{E}\left[\varphi(X_{T})\int_{0}^{T}(\sigma_{s}(X_{s})^{-1}J_{s\mid 0}\alpha_{s}^{\prime})^{\top}\mathrm{d}B_{s}\right](\alpha_{T}-\alpha_{0})^{-1}.$
Proposition 2.3 can now be proved using $\varphi=\delta_{x}$ in Theorem D.1. Intuitively, this corresponds to approximating the Dirac delta distribution by a sequence of peaked Gaussians $\varphi_{n}$, and then taking the limit. On a technical level, extending Theorem D.1 is supported by the framework of Watanabe distributions, see Watanabe (1987, Section 2).
Proof of Proposition 2.3.
Recall that the main ideas of the proof have already been outlined in the main text. Applying Theorem D.1 with $\varphi=\delta_{x_{T}}$, for fixed $x_{T}\in\mathbb{R}^{n}$, leads to
$\nabla\log p_{T\mid s}(x_{T}\mid x_{s})$ $=~{}\frac{1}{p_{T\mid s}(x_{T}\mid x_{s})}\nabla_{x_{s}}\mathbb{E}[\delta_{x_{T}}(X_{T})\mid X_{s}=x_{s}]$
$=~{}\mathbb{E}\left[\delta_{x_{T}}(X_{T})\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}\mid X_{s}=x_{s}\right]\frac{A_{T\mid s}^{-1}}{p_{T\mid t}(X_{T}\mid x_{s})}$
$=~{}\mathbb{E}\left[\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}\mid X_{s}=x_{s},X_{T}=x_{T}\right]A_{T\mid s}^{-1}.$
The final result of Proposition 2.3 follows transposing $\alpha$ and transposing the last equation; this is only a cosmetical change so that the result is more in line with algorithmic implementations.
Now, using Proposition D.2 for $t=T$ we get that
$\nabla\log p(Y=y|X_{s}=x_{s})$ $=\mathbb{E}[\nabla_{X_{t}}\log p_{t}(X_{T}|X_{s})|Y,X_{s}]$
$=\mathbb{E}\left[\mathbb{E}\left[\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}\mid X_{s},X_{T}\right]|Y,X_{s}\right]$
$=\mathbb{E}\left[\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}|Y,X_{s}\right]$
Proposition D.2.
Let $s\geqslant t$, then we have that
$\nabla_{X_{t}}\log p(Y|X_{s})=\mathbb{E}[\nabla_{X_{t}}\log p_{t}(X_{t}|X_{s})|Y,X_{s}].$
Proof.
We have
$\nabla_{X_{t}}\log p(Y|X_{t})$
$=$ $\frac{1}{p(Y|X_{t})}\nabla_{X_{t}}\int q(Y|X_{T}=x_{T})p(X_{T}=x_{T}|X_{t})\mathrm{d}x_{T}$
$=$ $\frac{1}{p(Y|X_{t})}\nabla_{X_{t}}\int q(Y|X_{T}=x_{T})p(X_{T}=x_{T}|X_{s}=x_{s})p(X_{s}=x_{s}|X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$=$ $\frac{1}{p(Y|X_{t})}\int q(Y|X_{T}=x_{T})p(X_{T}=x_{T}|X_{s}=x_{s})p(X_{s}=x_{s}|X_{t})\nabla_{X_{t}}\log p(X_{s}=x_{s}|X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$=$ $\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}|X_{t})|X_{t},Y].$
D.2 Equivalence to KL-Loss
Lemma D.3.
Letting $\mathcal{L}_{\operatorname{local}}^{t,y}(u_{t})$ be defined as in 16. Then it can be interpreted in terms of an amortized Kullback-Leibler divergence between measures on path space as follows:
$\int_{0}^{T}\mathcal{L}^{t}_{\mathrm{local}}(u_{t})\,\mathrm{d}t=\mathbb{E}\left[\mathrm{KL}(\mathbb{P}^{y}\mid\mathbb{P}^{u})\right]+C$
Proof.
We have that
$\mathcal{L}_{\operatorname{local}}^{t,y}(u_{t}):=$ $~{}\mathbb{E}[\|u(X_{t})-\mathcal{S}_{t}\|^{2}\mid Y=y]$
$=$ $~{}\mathbb{E}[\|u(X_{t})-(\mathcal{S}_{t}-\mathbb{E}[\mathcal{S}_{t}\mid X_{t},Y]+\mathbb{E}[\mathcal{S}_{t}\mid X_{t},Y])\|^{2}\mid Y=y]$
$=$ $~{}\mathbb{E}[\|u(X_{t})-\mathbb{E}[\mathcal{S}_{t}\mid X_{t},Y]\|^{2}\mid Y=y]+\mathbb{E}[\|\mathcal{S}_{t}-\mathbb{E}[\mathcal{S}_{t}\mid X_{t},Y]\|^{2}\mid Y=y]$
$=$ $~{}\mathbb{E}[\|u(X_{t})-\nabla\log p_{t}(Y=y\mid X_{t})\|^{2}\mid Y=y]+\mathbb{E}[\|\mathcal{S}_{t}-\mathbb{E}[\mathcal{S}_{t}\mid X_{t},Y]\|^{2}\mid Y=y]$
where we used that $\mathbb{E}[\mathcal{S}_{t}\mid X_{t},X_{T}]$ is an $L^{2}$ orthogonal projection of $\mathcal{S}_{s}$ onto the set of $\sigma(X_{t},X_{T})$ measurable random variables which $u(X_{t})$ is an element of. Integrating over this we obtain
$\int_{0}^{T}\mathcal{L}_{\operatorname{local}}^{t,y}(u_{t})\mathrm{d}t=\operatorname{KL}(\mathbb{P}^{y}\mid\mathbb{P}^{u})+\int_{0}^{T}\mathbb{E}[\|\mathcal{S}_{t}-\mathbb{E}[\mathcal{S}_{t}\mid X_{t},X_{T}]\|^{2}\mid X_{T}=x_{T}]\mathrm{d}t,$
where the second term is independent of $u$. Hence, assuming that the second term is integrable, the result follows by taking expectations on both sides. However, note that they have different finite-sample properties. ∎
D.3 Reparametrisation Method
The pathwise reparameterisation trick has been introduced by Domingo-Enrich et al. (2024b) in order to derive the stochastic optimal control matching loss. It is the following result:
Lemma D.4 ((Domingo-Enrich et al., 2024b), Prop. 1).
Let $X^{x}=(X^{x}_{t})$ be the solution of the SDE $\mathrm{d}X^{x}_{t}=b(X^{x}_{t},t)\,\mathrm{d}t+\sigma(t)\,\mathrm{d}B_{t}$ with initial condition $X^{x}_{0}=x$. Assume that $f:\mathbb{R}^{d}\times[0,T]\to\mathbb{R}$ and $g:\mathbb{R}^{d}\to\mathbb{R}$ are differentiable. For each $t\in[0,T]$, let $M_{t}:[t,T]\to\mathbb{R}^{d\times d}$ be an arbitrary continuously differentiable function matrix-valued function such that $M_{t}(t)=\mathrm{Id}$. We have that
$\begin{split}&\nabla_{x}\mathbb{E}\bigg{[}\exp\bigg{(}-\int_{0}^{T}f(X^{x}_{s},s)\,\mathrm{d}s-g(X^{x}_{T})\bigg{)}\bigg{]}\\ &~{}~{}~{}=\mathbb{E}\bigg{[}\bigg{(}-\int_{0}^{T}M_{s}\nabla_{x}f(X^{x}_{s},s)\,\mathrm{d}s-M_{T}\nabla g(X^{x}_{T})+\int_{0}^{T}(M_{s}\nabla_{x}b(X^{x}_{s},s)-M^{\prime}_{s})(\sigma^{-1})^{\top}(s)\mathrm{d}B_{s}\bigg{)}\\ &\qquad\qquad\times\exp\bigg{(}-\int_{0}^{T}f(X^{x}_{s},s)\,\mathrm{d}s-g(X^{x}_{T})\bigg{)}\bigg{]}.\end{split}$ (48)
Looking at the proof of this result in Domingo-Enrich et al. (2024b, Subsec. C.2), we observe that when $g$ is not differentiable, if we impose that $M_{T}=0$, then LABEL:eq:cond_exp_rewritten still holds. If we additionally set $f\equiv 0$ and $b$ time-independent, we obtain
$\nabla_{x}\mathbb{E}\big{[}\exp\big{(}-g(X^{x}_{T})\big{)}\big{]}=\mathbb{E}\bigg{[}\bigg{(}\int_{0}^{T}(M_{s}\nabla_{x}b(X^{x}_{s},s)-M^{\prime}_{s})(\sigma^{-1})^{\top}(s)\mathrm{d}B_{s}\bigg{)}\exp\big{(}-g(X^{x}_{T})\big{)}\bigg{]}.$ (49)
Next, we prove that relying on our approach, we can recover and generalise this result to compute $\nabla_{x}\mathbb{E}\big{[}\varphi(X^{x}_{T})\big{]}$ for $\varphi$ that are not necessarily strictly positive or differentiable:
Proof.
We apply Theorem D.1 for $\alpha_{s}=J_{s\mid 0}^{-1}M_{s}$ and observe that
$\sigma^{-1}_{s}J_{s\mid 0}\alpha_{s}^{\prime}=\sigma^{-1}_{s}(J_{s\mid 0}\alpha_{s})^{\prime}-\sigma^{-1}_{s}(J_{s\mid 0})^{\prime}\alpha_{s}=\sigma^{-1}_{s}M_{s}^{\prime}-\sigma^{-1}\nabla b(X_{s})J_{s\mid 0}\alpha_{s}=\sigma^{-1}_{s}(M_{s}^{\prime}-\nabla b(X_{s})M_{s}).$
Therefore
$\mathbb{E}[\nabla_{x}\varphi(X^{x}_{T})]$ $=\mathbb{E}\left[\varphi(X_{T})\int_{0}^{T}(\sigma^{-1}_{s}J_{s\mid 0}\alpha_{s}^{\prime})^{\top}\mathrm{d}B_{s}\right](\alpha_{T})^{-1}$
$=\mathbb{E}\left[\varphi(X_{T})\int_{0}^{T}(\sigma^{-1}_{s}(M_{s}^{\prime}-\nabla b(X_{s})M_{s}))^{\top}\mathrm{d}B_{s}\right](\alpha_{T}-\alpha_{0})^{-1}.$
Since $M_{s}$ is chosen in such a way that $M_{0}=\operatorname{Id}$ and $M_{T}=0$, we have $\alpha_{T}-\alpha_{0}=-\operatorname{Id}$, and the result follows. ∎
D.4 Gaussian Approximations
We will prove more general statements. Indeed, Lemma 3.3 still holds when we replace $X_{T}$ by a general observation $Y\sim q(X_{T},\cdot)$. To that end we first prove that the score is a martingale, when conditioned on the observation $Y$:
Lemma D.5.
Let $s\geqslant t$, then we have that
$\nabla_{X_{t}}\log p(Y\mid X_{t})=\mathbb{E}[\nabla_{X_{t}}\log p_{t}(X_{s}\mid X_{t})\mid Y,X_{t}].$ (50)
Proof.
We have
$\nabla_{x_{t}}\log p(Y\mid X_{t})$
$\qquad=\frac{1}{p(Y\mid X_{t})}\nabla_{x_{t}}\int q(Y\mid X_{T}=x_{T})p(X_{T}=x_{T}\mid X_{t})\mathrm{d}x_{T}$
$\qquad=\frac{1}{p(Y\mid X_{t})}\nabla_{x_{t}}\int q(Y\mid X_{T}=x_{T})p(X_{T}=x_{T}\mid X_{s}=x_{s})p(X_{s}=x_{s}\mid X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$\qquad=\frac{1}{p(Y\mid X_{t})}\nabla_{x_{t}}\int q(Y\mid X_{T}=x_{T})p(X_{T}=x_{T}\mid X_{s}=x_{s})p(X_{s}=x_{s}\mid X_{t})\nabla_{X_{t}}\log p(X_{s}=x_{s}\mid X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$\qquad=\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y].$
Now we use the above to show the equivalence of the two losses:
Lemma D.6.
(Lemma 3.3) It holds that
$\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p(Y\mid X_{t})\|^{2}]=C+\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\|^{2}].$ (51)
Proof.
We have that
$\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p(Y\mid X_{t})\|^{2}]=\mathbb{E}[\|u_{t}(X_{t},Y)-\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]$
by Lemma D.5. Since the conditional expectation $\mathbb{E}[\cdot\mid X_{t},Y]$ is a orthogonal projection onto the subspace of $\sigma(X_{t},Y)$-measurable random variables in $L^{2}$, and $s(t,X_{t},Y)$ is an element of that subspace, we have that
$\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\|^{2}]$
$~{}~{}=\mathbb{E}[\|u_{t}(X_{t},Y)-\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]+\mathbb{E}[\|\nabla_{X_{t}}\log p(X_{s}\mid X_{t})-\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]$
$~{}~{}=\mathbb{E}[\|u_{t}(X_{t},Y)-\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]+\mathbb{E}[\|\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\|^{2}]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\mathbb{E}[\|[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}].$
This proves the statement. ∎
Finally, we prove
Lemma D.7 (Lemma 3.4).
Approximating $p(X_{t+\delta t}\mid X_{t})$ by a Gaussian (24a) and regressing against its score is equivalent to choosing $\alpha^{\prime}_{s}=1_{[t,t+\delta t]}$ and approximating the stochastic integral in $\mathcal{S}_{s}$ (8) as
$\int_{t}^{t+\delta t}J_{s\mid t}\mathrm{d}B_{s}\approx J_{t+\delta t\mid t}(B_{t+\delta t}-B_{t}),$ (52)
and furthermore approximating $J_{t+\delta t\mid t}$ by an Euler-Maryuama step on (7):
$J_{t+\delta t\mid t}\approx\operatorname{Id}+\delta t\nabla b(t,X_{t}).$ (53)
Proof.
Since we approximate
$p(X_{t+\delta t}\mid X_{t})\approx\mathcal{N}(X_{t}+\delta tb_{t}(X_{t}),\delta t),$
we have that
$\nabla\log p(X_{t+\delta t}\mid X_{t})=(\text{Id}+\delta t\nabla b(t,X_{t}))(X_{t+\delta t}-(X_{t}+\delta tb(X_{t})))=\frac{1}{\delta t}(\text{Id}+\delta t\nabla b(t,X_{t}))(B_{t+\delta t}-B_{t}).$ (54)
Since $\delta=\alpha_{t+\delta t}-\alpha_{t}$, plugging the normalization $\frac{1}{\alpha_{t+\delta t}-\alpha_{t}}=\frac{1}{\delta}$ into $\mathcal{S}_{s}$ (8) shows that the two expressions (8) and (54) are the same. ∎
Appendix E Gaussian Analysis of the Variance
In this Section, we prove Lemma 3.1. First we determine the variance of the Monte Carlo estimator:
Lemma E.1.
Set $T=1$, $n=1$, $b=0$ and $\sigma=1$, i.e., $X_{t}$ is a one-dimensional Brownian motion conditioned on $X_{0}=x_{0}$ and $X_{1}=x_{1}$. Then the variance of the Monte-Carlo estimator
$\nabla\log p_{1\mid 0}(B_{1}=x\mid B_{0}=x_{0})\approx\int_{0}^{1}\alpha_{t}^{\prime}J^{\top}_{t\mid 0}(\sigma_{t}^{-1})^{\top}\mathrm{d}B_{t}$
is given by
$\frac{1}{\alpha_{1}-\alpha_{0}}\!\left(\int_{0}^{1}\left({\alpha^{\prime}_{t}}\right)^{2}\mathrm{d}t+\!\int_{0}^{1}\frac{(\alpha_{1}-\alpha_{t})^{2}}{(1-t)^{2}}\mathrm{d}t+(x_{1}-x_{0})^{2}\!\right)\!,$ (55)
assuming that $\frac{\alpha_{1}-\alpha_{s}}{\sqrt{1-s}}\rightarrow 0$ as $s\rightarrow 1$.
Proof.
Without loss of generality, we assume that $\alpha_{s}^{\prime}$ integrates to $1$, and we use the notation $x\equiv x_{0}$ and $y\equiv x_{T}$. We need the following preliminary facts,
$\mathbb{E}[B_{s}\mid B_{1}=y,B_{0}=x]=\mathbb{E}[W_{s}-sW_{1}+(1-s)x+sy]=(1-s)x+sy$
$\mathbb{E}[\nabla\log p(B_{1}=x\mid B_{s})\mid B_{1}=y,B_{0}=x]=\mathbb{E}\left[\frac{x-B_{s}}{1-s}\mid B_{1}=y,B_{0}=x\right]=(y-x),$
where $W_{s}$ is a standard (unconditioned) Brownian motion. The expectation of the Monte Carlo estimator can then be computed as
$\mathbb{E}\left[\int_{0}^{1}\sigma_{s}^{-1}J_{s\mid 0}\alpha_{s}^{\prime}\mathrm{d}B_{s}\mid B_{1}=y,B_{0}=x\right]$ $=\mathbb{E}\left[\int_{0}^{1}\alpha_{s}^{\prime}\mathrm{d}B_{s}\mid B_{1}=y,B_{0}=x\right]$
$=\mathbb{E}\left[\int_{0}^{1}\alpha_{s}^{\prime}\mathrm{d}(W_{s}+\nabla\log p(B_{1}=x\mid B_{s})\mathrm{d}s)\mid B_{1}=y,B_{0}=x\right]$
$=\mathbb{E}\left[\int_{0}^{1}\alpha_{s}^{\prime}\nabla\log p(B_{1}=x\mid B_{s})\mathrm{d}s\mid B_{1}=y,B_{0}=x\right]$
$=\mathbb{E}\left[\int_{0}^{1}\alpha_{s}^{\prime}\nabla\log p(B_{1}=x\mid B_{s})\mathrm{d}s\right]$
$=\int_{0}^{1}(y-x)\alpha_{s}^{\prime}\mathrm{d}s=(y-x).$
For the square of the estimator we obtain
$\mathbb{E}\left[\left(\int_{0}^{1}\sigma_{s}^{-1}J_{s\mid 0}\alpha_{s}^{\prime}\mathrm{d}B_{s}\right)^{2}\right]$ $=\mathbb{E}\left[\left(\int_{0}^{1}\alpha_{s}^{\prime}(\mathrm{d}W_{s}+\nabla_{B_{s}}\log p(B_{1}=x\mid B_{s})\mathrm{d}s)\right)^{2}\mid B_{1}=y\right]$
$=\mathbb{E}\left[\left(\int_{0}^{1}\alpha_{s}^{\prime}\mathrm{d}W_{s}\right)^{2}\right]+\mathbb{E}\left[\int_{0}^{1}\alpha_{s}^{\prime}\mathrm{d}W_{s}\int\alpha_{s}^{\prime}(y-x)\mathrm{d}s\right]+\mathbb{E}\left[\left(\int_{0}^{1}\alpha_{s}^{\prime}\frac{y-B_{s}}{1-s}\mathrm{d}s\right)^{2}\right]$
$=\int_{0}^{1}(\alpha_{s}^{\prime})^{2}\mathrm{d}s+0+\mathbb{E}\left[\left(\int\alpha_{s}^{\prime}\frac{y-B_{s}}{1-s}\mathrm{d}s\right)^{2}\right].$
Moreover,
$\mathbb{E}\left[\left(\int\alpha_{s}^{\prime}\frac{y-B_{s}}{1-s}\mathrm{d}s\right)^{2}\right]$ $=\mathbb{E}\left[\int_{0}^{1}\int_{0}^{1}\alpha_{s}^{\prime}\frac{B_{s}-y}{1-s}\alpha_{t}^{\prime}\frac{B_{t}-y}{1-t}\mathrm{d}s\,\mathrm{d}t\right]$
$=\int_{0}^{1}\int_{0}^{1}\left(\alpha_{s}^{\prime}\alpha_{t}^{\prime}\frac{\min(s,t)-st}{(1-s)(1-t)}+\frac{\alpha_{s}^{\prime}\alpha_{t}^{\prime}\mathbb{E}[B_{s}-y]\mathbb{E}[B_{t}-y]}{(1-s)(1-t)}\right)\mathrm{d}s\,\mathrm{d}t$ (56)
The first term in 56 can simplified,
$\int_{0}^{1}\int_{0}^{1}\alpha_{s}^{\prime}\alpha_{t}^{\prime}\frac{\min(s,t)-st}{(1-s)(1-t)}$ $=\int_{0}^{1}\int_{0}^{t}\alpha_{s}^{\prime}\alpha_{t}^{\prime}\frac{s-st}{(1-s)(1-t)}\mathrm{d}s\,\mathrm{d}t+\int_{0}^{1}\int_{t}^{1}\alpha_{s}^{\prime}\alpha_{t}^{\prime}\frac{t-st}{(1-s)(1-t)}\mathrm{d}s\,\,\mathrm{d}t$
$=\int_{0}^{1}\int_{0}^{t}\alpha_{s}^{\prime}\alpha_{t}^{\prime}\frac{s(1-t)}{(1-s)(1-t)}\mathrm{d}s\,\mathrm{d}t+\int_{0}^{1}\int_{t}^{1}\alpha_{s}^{\prime}\alpha_{t}^{\prime}\frac{t(1-s)}{(1-s)(1-t)}\mathrm{d}s\,\mathrm{d}t$
$=\int_{0}^{1}\int_{0}^{t}\alpha_{s}^{\prime}\alpha_{t}^{\prime}\frac{s}{(1-s)}\mathrm{d}s\,\mathrm{d}t+\int_{0}^{1}\int_{t}^{1}\alpha_{s}^{\prime}\alpha_{t}^{\prime}\frac{t}{(1-t)}\mathrm{d}s\,\mathrm{d}t$
$=\int_{0}^{1}\alpha_{s}^{\prime}\frac{s}{(1-s)}\int_{s}^{1}\alpha_{t}^{\prime}\mathrm{d}t\,\mathrm{d}s+\int_{0}^{1}\alpha_{t}^{\prime}\frac{t}{(1-t)}\int_{t}^{1}\alpha_{s}^{\prime}\mathrm{d}s\,\mathrm{d}t$
$=\int_{0}^{1}\alpha_{s}^{\prime}\frac{s}{(1-s)}(\alpha_{1}-\alpha_{s})\,\mathrm{d}s$
$=2\int_{0}^{1}\frac{s}{1-s}u(s)^{\prime}\,\mathrm{d}s,$
where we have defined $u(s)\coloneqq\frac{1}{2}(\alpha_{1}-\alpha_{s})^{2}.$ Now we see that
$2\int_{0}^{1}\frac{s}{1-s}u(s)^{\prime}\,\mathrm{d}s$ $=-2{\left[u(s)\frac{s}{1-s}\right]_{0}^{1}}+2\int_{0}^{1}\left(\frac{s}{1-s}\right)^{\prime}u(s)\,\mathrm{d}s$
$=0+2\int_{0}^{1}\frac{1}{(1-s)^{2}}u(s)\,\mathrm{d}s$
$=\int_{0}^{1}\frac{(\alpha_{1}-\alpha_{s})^{2}}{(1-s)^{2}}\mathrm{d}s,$
where we have used the fact that $\alpha_{1}-\alpha_{s}=o\left(\sqrt{1-s}\right)$ by assumption. The second term in (56) simplifies as follows:
$\int_{0}^{1}\int_{0}^{1}\frac{\alpha_{s}^{\prime}\alpha_{t}^{\prime}\mathbb{E}[B_{s}-y]\mathbb{E}[B_{t}-y]}{(1-s)(1-t)}\mathrm{d}s\,\mathrm{d}t$ $=\int_{0}^{1}\int_{0}^{1}\frac{\alpha_{s}^{\prime}\alpha_{t}^{\prime}(x-y)((1-s))(x-y)(1-t)}{(1-s)(1-t)}\mathrm{d}s\,\mathrm{d}t$
$=(x-y)^{2}\int_{0}^{1}\int_{0}^{1}\alpha_{s}^{\prime}\alpha_{t}^{\prime}\frac{}{}\mathrm{d}s\,\mathrm{d}t=(x-y)^{2},$
so that collecting all the terms yields the claimed result. ∎
In the following determine the optimal choice of $\alpha$ in the simplified setting:
Lemma E.2.
Assume the setting from Lemma 3.1. Then the optimal $\alpha$ has derivative
$\alpha_{s}^{\prime}=\left(\frac{1}{2}\left(1+\sqrt{5}\right)\right)(1-s)^{\frac{1}{2}\left(-1+\sqrt{5}\right)}.$
Proof.
We assume without loss of generality that $\alpha_{1}=1$ and $\alpha_{0}=0$. We need to optimise the following term,
$\int_{0}^{1}\left({\alpha^{\prime}_{s}}\right)^{2}\mathrm{d}s+\int_{0}^{1}\frac{(\alpha_{1}-\alpha_{s})^{2}}{(1-s)^{2}}\,\mathrm{d}s=\int_{0}^{1}\left({\beta_{1-s}}^{\prime}\right)^{2}+\frac{(\beta_{1-s})^{2}}{(1-s)^{2}}\,\mathrm{d}s=-\int_{0}^{1}\left({\beta_{s}}^{\prime}\right)^{2}+\frac{(\beta_{s})^{2}}{s^{2}}\,\mathrm{d}s,$
where we have set $\beta_{s}\coloneqq\alpha_{1}-\alpha_{1-s}.$ We define
$F(\beta):=\int_{0}^{1}\left({\beta_{s}}^{\prime}\right)^{2}+\frac{(\beta_{s})^{2}}{s^{2}}\mathrm{d}s,$
since the last term in 19 is independent of $\beta_{s}$. Assuming that $\alpha$ is a minimizer of $F$, it follows that $\frac{\mathrm{d}}{\mathrm{d}\varepsilon}F(\beta+\varepsilon\varphi)=0$ for any $\varphi$ such that the $\alpha$ defined through $\beta+\varepsilon\varphi$ satisfies the assumptions of Lemma 3.1. In particular, this condition has to be satisfied for any weakly differentiable $\varphi$ such that $\varphi_{0}=\varphi_{1}=1$. Given that, we calculate
$\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\int_{0}^{1}\left(\left({\beta_{s}}^{\prime}+\varepsilon\varphi\right)^{2}+\frac{(\beta_{s}+\varepsilon\varphi)^{2}}{s^{2}}\right)\mathrm{d}s=2\int_{0}^{1}\left(\varphi^{\prime}\left({\beta_{s}}^{\prime}+\varepsilon\varphi^{\prime}\right)+\frac{\varphi(\beta_{s}+\varepsilon\varphi)}{s^{2}}\right)\mathrm{d}s.$
Therefore, we arrive at
$0\mathop{=}\limits^{!}\int_{0}^{1}\varphi^{\prime}{{\beta_{s}}^{\prime}}+\frac{\varphi\beta_{s}}{s^{2}}\mathrm{d}s=\int_{0}^{1}{{{-\varphi\beta_{s}}^{\prime}}^{\prime}}+\frac{\varphi\beta_{s}}{s^{2}}\mathrm{d}s.$
Since this equation needs to hold for any continuous $\varphi$ with 00-boundary conditions it follows that
$\beta_{s}^{\prime\prime}=\frac{\beta_{s}}{s^{2}}.$
Solving this with the boundary conditions $\beta_{0}=0$ and $\beta_{1}=1$ leads to
$\beta_{s}$ $=s^{\frac{1}{2}\left(1+\sqrt{5}\right)},$
$\beta_{s}^{\prime}$ $=\left(\frac{1}{2}\left(1+\sqrt{5}\right)\right)s^{\frac{1}{2}\left(-1+\sqrt{5}\right)},$
as claimed. ∎